OpenStudy (anonymous):

Where do the lines modeled by the equations y= -7/8x + 1/4 and y= -3/4x + 1/2 intersect? A) (-4, 2) B) (-2,2) C) (2,-2) D) (4,-2)

OpenStudy (johnweldon1993):

Since 'y' equals both of these equations....you can set them equal to each other...and solve for 'x' \[\large -\frac{7}{8}x + \frac{1}{4} = -\frac{3}{4}x + \frac{1}{2}\]

OpenStudy (anonymous):

Working of of what John said, you can plug the x-value you find into either of the equations to find the y-value of the intersection.

OpenStudy (anonymous):

So how do you find where the line intersect?

OpenStudy (anonymous):

are you there ?

OpenStudy (texaschic101):

-7/8x + 1/4 = -3/4x + 1/2 -- multiply by 8 -7x + 2 = -6x + 4 -7x + 6x = 4 - 2 -x = 2 x = -2 now sub -2 in for x in either equation to find y y = -3/4x + 1/2 y = -3/4(-2) + 1/2 y = 6/4 + 1/2 y = 6/4 + 2/4 y = 8/4 y = 2 check. y = -7/8x + 1/4 2 = -7/8(-2) + 1/4 2 = 14/8 + 1/4 2 = 14/8 + 2/8 2 = 16/8 2 = 2 (correct) the lines intersect at (-2,2)